OpenStudy (anonymous):

6 x (5+7) /4-7x2+3x3

OpenStudy (aishalicious1):

What am I looking at?

OpenStudy (okdutchman7):

What do you already know how to do?

OpenStudy (okdutchman7):

What part are you stuck on?

OpenStudy (aishalicious1):

Explain what you need help with

OpenStudy (okdutchman7):

@cmwheeler03

OpenStudy (aishalicious1):

Hellooooooo?

OpenStudy (boothepest1):

6x(5+7)/4-7x2+3x3 = 6x(12)/4-14+9 =18-14+9 =13

OpenStudy (boothepest1):

@Cmwheeler03 Hope that helps!

MsBrains (ms-brains):

Hi! Welcome to open study! What @Boothepest1 said seems correct. :) 6x(5+7)/4-7x2+3x3 = 6x(12)/4-14+9 =18-14+9 =13